OpenStudy (anonymous):

James works at a video store. He makes $6.20 per hour and works 30 hours each week. How much money will James make in a 4 week period? $ ___________

OpenStudy (anonymous):

17856 is the answer i have gotten but sure that it is correct

OpenStudy (anonymous):

$ 744

OpenStudy (anonymous):

Each hour James makes $6.20, so 30 hours he makes 6.20*30=186 dollars. He makes 186*4=744 dollars in a period of 4 weeks.

OpenStudy (anonymous):

Just multiply his wage (6.20) by his hours (30) to get 186. Then multiply 186 by the number of weeks in a month (4) to get $744. \[4(6.20*30)=744\]

OpenStudy (anonymous):

Good, Just_Al!

OpenStudy (anonymous):

oh i kinda figured by each hour in a day ehich is 24 and multiplied it but 2 and then multiplied the answers together 6.20 times 30 times etc

OpenStudy (anonymous):

That wouldn't be how it works, Starr. He is ONLY working 30 hours during a week, which means you only multiple 6.20 by 30.